OpenStudy (anonymous):

Suppose the average annual expense for a random sample of 12 stock funds is normally distributed with mean of 1.63 percent with a standard deviation of .31 percent. Calculate a 95 percent confidence interval for the mean annual expense charged by all stock funds.

OpenStudy (kropot72):

The confidence interval is found by plugging the given values into the following: \[\large (\bar{x}-1.96\frac{\sigma}{\sqrt{n}},\ \bar{x}+\frac{\sigma}{\sqrt{n})})\]

OpenStudy (kropot72):

Giving: \[\large CI=(1.63-1.96\frac{0.31}{\sqrt{12}},\ 1.63+1.96\frac{0.31}{\sqrt{12}})=you\ can\ calculate\]

OpenStudy (kropot72):

@work12345 Are you there?

OpenStudy (anonymous):

@kropot72 thanks a lot for your help! Sorry about that, I left my computer after waiting awhile on this question.

OpenStudy (kropot72):

You're welcome :)